OpenStudy (anonymous):

f(x) = 3x^2 -2x, find f(a+1)

OpenStudy (anonymous):

Just plug in a + 1 in for x. \[f(x) = 3x^2 - 2x\]\[f(a + 1) = 3(a + 1)^2 - 2(a + 1) \implies ?\]